# Geometrically Thick Disks around Kerr Black Holes in a Swirling Universe

### v1

@@ 4 Circular Orbits and Disk solutions
Using the same disk parameters besides the Kerr parameter reveals, that prograde disk solutions are in their general morphology largely unaffected by the Kerr parameter. All disk solutions are asymmetrical to the equatorial plane and their spatial distribution is similar, regardless of the Kerr parameter. They extend mainly into the upper hemisphere with only a small region extending into the lower hemisphere. This effect originates from the $\mathcal{Z}_{2}$ symmetry of the swirling background and increases with increasing $j$. We follow, that even in the vicinity of the KBHSU, the Kerr rotation gets dominated by the swirling rotation albeit $j$ being small compared to $a$. However, differences regarding the physical disk properties are noticeable. For the same disk parameters, the effective potential for the slower Kerr rotation has a deeper minimum and therefore a higher density at the center of the disk compared to the faster rotation. The disk solution for $a=0.2$ is of Type 1a, where the disk center is enclosed by the equi-potential surface of the outer cusp, which itself is surrounded by the equi-potential surface of the inner cusp (Fig. 10 (a)). The $a=0.5$ and $a=0.9$ disk solutions are of Type 2, with only an outer cusp. Since they do not possess an inner cusp, no accretion process without further inner-physical disk effects is possible. For faster rotating KBHSUs, prograde disks solutions exist for higher $j$ and for lower $\ell_{0}$, leading to more compact disks, as shown in Fig. 11.
@@
In this Type 1b solution, the center region is highly compact, most of the matter is located close the density maximum (minimum of $\mathcal{W}$) and the disk center is located closely to the KBHSU center. The effective potential decreases monotonically outwards the outer cusp also on a greater scale. We follow, that with increasing $a$ and $j$ and decreasing $\ell_{0}$ disk solutions are more localized and compact. The verticality of the disk structure is mainly affected by the swirling parameter. Since with increasing $a$ disk solutions are possible for larger values of $j$, solutions with a greater verticality are possible for faster rotating KBSHU.
In the following Fig. 12 examples of retrograde disk solutions are presented for various values of $a$, $j$ and $\ell_{0}$. Due to the more limited disk solution space for retrograde disks, solutions with the same $\ell_{0}$ and $j$ value do not exist for the picked Kerr parameters ($a\in\{0.2,0.5,0.9\}$). A comparison for different Kerr parameters but same swirling and disk parameters, like done for the prograde disks, is therefore not possible. Thus, we picked exemplary solutions with different parameters to showcase the main properties of retrograde disks.
@@
Retrograde disks are similar to prograde disks regarding their asymmetry with reference to the equatorial plane. In contrast to prograde disks, they extend mainly into the lower hemisphere with only a small part extending into the upper hemisphere. This corresponds to the spatial distribution of the retrograde circular orbits, which originates from the change of the swirling rotation in the lower hemisphere, and with increasing $j$ this asymmetry and verticality of the disk increases. The Figs. 12 (a) and (b) present retrograde Type 1b disks, for which the same properties apply as explained for the prograde case. Fig. 12 (c) illustrates a retrograde Type 1c disk, where the effective potential at the inner and outer cusp has the same value. This value is also closely located to the potential minimum. The potential difference between the disk center and inner/outer cusp is miniscule, which causes the enclosed region between inner and outer cusp to appear as vanishing. The only stable region of the disk lies inside this small narrow region close to the disk center. Regarding disk dynamics this solution could therefore be highly unstable. The following Fig. 13 presents a retrograde disk solution for the fast rotating KBSHU with $a=0.9$.
@@
This presented solution is of Type 1a, with the equi-potential surface of the outer cusp enclosing the disk center. The main disk properties are similar as explained for the prograde Type 1a disks. Nevertheless, for faster rotating KBSHU the retrograde disk center moves outwards and the disks are less compact compared to other disk solutions. This is also reflected by the large scale structure (Fig. 13 (b)), where the horizontal extension is greater compared to the prograde case for $a=0.9$, due to the smaller swirling parameter $j$. A greater verticality would correlate with a higher value of the Swirling parameter, however the retrograde disk solution space regarding the Swirling parameter is very limited in case of faster rotating KBSHU, since for higher $j$ no stable circular orbits and therefore no disk solutions exist.
@@ 5 Conclusion
In this work, we investigated qualitatively geometrically thick disks around Kerr black holes in a swirling universe. We classified different KBHSU solutions by their Kerr spin parameter $a$ and varied for fixed $a$ the swirling parameter $j$, to analyze the influence of the swirling background and the effects of the spin-spin interaction between black hole and background. For an exemplary analysis regarding the circular orbits and disk solutions, we chose the Kerr parameters $a=0.2$, $a=0.5$, and $a=0.9$, for slow, medium, and fast rotating black holes, respectively. The swirling parameter was varied between $j=10^{-5}$ and $5\cdot 10^{-4}$, as for higher values of $j$ no disk solutions exist. We analyzed prograde orbits and disks, as well as retrograde orbits and disks. In both cases, the properties of the orbits and disks are highly affected by the swirling parameter, as circular orbits deviate from the equatorial plane with increasing $j$. Prograde (retrograde) orbits get shifted into the upper (lower) hemisphere, where they are comoving with the swirling background and the verticality of the spatial distribution increases with $j$ for all possible values of $a$. The differences regarding the spatial distribution for the different Kerr parameters are negligible. The Keplerian specific angular momentum distribution develops in all cases an additional outer extremum, which marks an outer marginally stable orbit. The range of stable orbits is therefore bounded by the inner and outer marginally stable orbit, which also severely limits the parameter range of possible disk solutions.
In the case of prograde orbits, we found a stabilizing effect of the spin-spin interaction, since for higher values of $a$ stable circular orbits were possible for larger $j$. Faster rotating black holes harbour therefore a broader parameter space for prograde disk solutions. For retrograde orbits, which are all located in the lower hemisphere, we found the opposite effect. In the lower hemisphere, the black hole is counter-rotating to the swirling background, which has a destabilizing effect on the circular orbits. Therefore with increasing $a$ the range of stable circular orbits decreases for larger $j$ and the parameter space for disk solutions shrinks.
@@
Due to the outer extremum of the Keplerian specific angular momentum all possible disk solutions possess an outer cusp. In dependence of the spacetime solutions and the chosen value of the disk momentum $\ell_{0}$ we found and classified four different possible disk types. Type 1 solutions are characterized by an inner cusp and an outer cusp, between which the disk center lies. Type 2 solutions only exist for prograde disks for the anaylzed values of $a$ and are characterized by the absence of an inner cusp, they only possess an outer cusp. The Type 1 solutions can be further subclassified regarding the potential values at the cusps. For Type 1a solutions, the effective potential has a higher value at the inner cusp than the outer cusp. This causes the equi-potential surfaces of the inner cusp to be open towards the center and towards the outer disk regions. The equi-potential surface of the outer cusp fully encloses the disk center region and is closed towards the KBHSU center but open towards the outer region. In the case of Type 1b solutions, the effective potential of the outer cusp has a higher value than the inner cusp. This causes the equi-potential surface of the outer cusp to be open towards the center and towards the outer region. The inner cusp potential surface now fully encloses the disk center, it is closed towards the outer region and open towards the KBHSU center. In the special case Type 1c the effective potential has the same value at the inner and outer cusp, the potential surface is fully closed between the inner and outer cusp and and it encloses the disk center, but is open towards the KBHSU center and towards the outer disk region. All mentioned solutions have a high complexity of disk dynamics due to the different conceivable accretion mechanisms regarding the inner as well as the outer cusp. However, direct accretion onto the KBHSU, with the absence of inner-physical disk effects, is only possible for Type 1 solutions as Type 2 do not possess an inner cusp and are closed towards the KBSHU center.
@@ Acknowledgements.
We would like to thank Betti Hartmann and Jutta Kunz for discussions. R.C. is grateful to CAPES for financial support under Grant No: 88887.371717/2019-00, and would like to thank the University of Oldenburg for hospitality.

### v2

@@ 4 Circular Orbits and Disk solutions
Using the same disk parameters besides the Kerr parameter, reveals that prograde disk solutions are in their general morphology largely unaffected by the Kerr parameter. All disk solutions are asymmetrical to the equatorial plane, and their spatial distribution is similar, regardless of the Kerr parameter. They extend mainly into the upper hemisphere with only a small region extending into the lower hemisphere. This effect originates from the $\mathcal{Z}_{2}$ symmetry of the swirling background and increases with increasing $j$. We follow, that even in the vicinity of the KBHSB, the Kerr rotation gets dominated by the swirling rotation, albeit $j$ being small compared to $a$. However, differences regarding the physical disk properties are noticeable. For the same disk parameters, the effective potential for the slower Kerr rotation has a deeper minimum and therefore a higher density at the center of the disk compared to the faster rotating black hole. The disk solution for $a=0.2$ is of Type 1a, where the disk center is enclosed by the equi-potential surface of the outer cusp, which itself is surrounded by the equi-potential surface of the inner cusp (Fig. 10 (a)). The $a=0.5$ and $a=0.9$ disk solutions are of Type 2, with only an outer cusp. Since they do not possess an inner cusp, no accretion process without further inner-physical disk effects is possible. For faster rotating KBHSBs, prograde disks solutions can exist for higher $j$ and smaller values of lower $\ell_{0}$, leading to more compact disks, as shown in Fig. 11.
@@
In this Type 1b solution above, the center region is highly compact, most of the matter is located close the density maximum (minimum of $\mathcal{W}$) and the disk center is located closely to the KBHSB center. The effective potential decreases monotonically outwards the outer cusp, also on a greater scale. We follow that with increasing $a$ and $j$ and decreasing $\ell_{0}$ disk solutions are more localized and compact. The verticality of the accretion structure is mainly affected by the swirling parameter. Since with increasing $a$ disk solutions are possible for larger values of $j$, solutions with a greater verticality are possible for faster rotating KBHSB.
In the following Fig. 12, examples of retrograde disk solutions are presented for various values of $a$, $j$ and $\ell_{0}$. Due to the more limited parameter space for retrograde accretion structures, solutions with the same $\ell_{0}$ and $j$ value do not exist for the picked Kerr parameter ($a\in\{0.2,0.5,0.9\}$). A comparison for different Kerr parameter but same swirling and disk parameters, like done for the prograde disks, is therefore not possible. Thus, we picked exemplary solutions with different parameters to showcase the main properties of retrograde disks.
@@
Retrograde disks are similar to prograde disks regarding their asymmetry with respect to the equatorial plane. In contrast to prograde disks, they extend mainly into the lower hemisphere with only a small part extending into the upper hemisphere. This corresponds to the spatial distribution of the retrograde circular orbits, which originates from the change of the swirling rotation in the lower hemisphere. With increasing $j$, the asymmetry and verticality of the accretion structure increases. The Figs. 12 (a) and (b) present retrograde Type 1b disks, for which the same properties apply as explained for the prograde case. The Fig. 12 (c) illustrates a retrograde Type 1c disk, where the effective potential at the inner and outer cusp has the same value. This value is also closely located to the potential minimum. The potential difference between the disk center and inner/outer cusp is minuscule, which causes the enclosed region between inner and outer cusp to appear as vanishing. The only stable region of the structure lies inside this small narrow region close to the disk center. Regarding disk dynamics this solution could therefore be highly unstable. The following Fig. 13 presents a retrograde disk solution for the fast rotating KBHSB with $a=0.9$.
@@
The presented solution above is of Type 1a, with the equi-potential surface of the outer cusp enclosing the disk center. The main disk properties are similar to the prograde Type 1a disks. Nevertheless, for faster rotating KBHSB, the retrograde disk center moves outwards and the disks are less compact compared to other disk solutions. This is also reflected by the large scale structure (Fig. 13 (b)), where the horizontal extension is greater compared to the prograde case for $a=0.9$ due to the smaller swirling parameter $j$. A greater verticality would correlate with a higher value of the Swirling parameter; however, the retrograde disk parameter space regarding the swirling parameter is very limited in the case of faster rotating KBHSB, since for higher $j$ no stable circular geodesics and therefore no disk solutions exist.
5 Comparison to Thick Disks around Schwarzschild Black Holes in a Swirling Background
Ahead of the conclusion, we want to conduct a brief comparison to thick disks around a Schwarzschild black hole in a swirling background, which were studied in [37], and point out the effects of the black hole rotation and their consequences. We identify the spin-spin interaction of the rotating black hole with the swirling rotation of the spacetime background as a major differentiating factor for possible disk solutions. The spin-spin interaction influences the stability properties of circular geodesics and by that, it also affects the parameter space for thick disk solutions. The Kerr spin parameter determines the critical Swirling parameter, $j_{c}=j_{c}(a)$, up to which stable circular geodesics exist in the spacetime. For prograde motion, $j_{c}$ is an increasing function of $a$, while for retrograde motion, it is a decreasing function of $a$. If the value of the swirling parameter is fixed, the black hole rotation determines the spectrum of specific angular momenta for the disk particles, for which solutions exist (Fig. 8). In the case of prograde motion, a higher value of the Kerr spin parameter leads to a greater range of possible values of the specific angular momentum, while for retrograde motion the opposite holds. However, the vertical position of circular geodesics is mostly unaffected by the Kerr spin parameter, hence it is mainly determined by the Swirling parameter. Thus, the spatial distribution of circular geodesics stays similar to the Schwarzschild case for all Kerr spin parameters (Fig. 7). Additionally, the positions of static orbits are mostly unaffected by the choice of the Kerr parameter (Fig. 6). When it comes to the disk morphology and matter distribution, the Kerr parameter has a smaller influence, the disk shape is mainly determined by the Swirling parameter. However, disk solutions for Kerr spin parameter near the Kerr limit are more vertically bent and less horizontally outreaching. This applies to prograde as well as retrograde disks. Finally, we conclude that the major differences resulting from the spin-spin interaction appear in the range of the parameter space for which stable circular geodesic and by that also accretion structures can exist.
@@ 6 Conclusion
In this work, we investigated qualitatively geometrically thick disks around Kerr black holes in a swirling background. We classified different KBHSB solutions by their Kerr spin parameter $a$ and varied for fixed $a$ the swirling parameter $j$ to analyze the influence of the swirling background and the effects of the spin-spin interaction between the black hole and the background. For an exemplary analysis regarding the circular orbits and disk solutions, we chose the Kerr parameters $a=0.2$, $a=0.5$, and $a=0.9$, for slow, medium, and fast rotating black holes, respectively. The swirling parameter was varied between $j=10^{-5}$ and $5\cdot 10^{-4}$, as for higher values of $j$ no stable circular orbits exist. Fruthermore, we have considered prograde and retrograde motion. In both cases, the properties of the orbits and disks are highly affected by the swirling parameter, as circular orbits deviate from the equatorial plane with increasing $j$. Prograde (retrograde) orbits get shifted into the upper (lower) hemisphere, where they are co-moving with the swirling background, and the verticality of the spatial distribution increases with $j$ for all possible values of $a$. The differences regarding the spatial distribution for the different Kerr parameters are negligible. The Keplerian specific angular momentum distribution develops in all cases an additional outer extremum, which marks an outer marginally stable orbit. The range of stable orbits is therefore bounded by the inner and outer marginally stable orbit, which also severely limits the parameter space of possible disk solutions.
In the case of prograde orbits, we found a stabilizing effect of the spin-spin interaction, since for higher values of $a$, stable circular orbits were possible for larger $j$. Faster rotating black holes therefore exhibit a broader parameter space for prograde disk solutions. For retrograde orbits, which are all located in the lower hemisphere, we found the opposite effect. In the lower hemisphere, the black hole is counter-rotating to the swirling background, which has a destabilizing effect on the circular orbits. Therefore, with increasing $a$, the range of stable circular orbits decreases for larger $j$ and the parameter space for disk solutions shrinks.
@@
Due to the outer extremum of the Keplerian specific angular momentum, all possible accretion structures possess an outer cusp. In dependence of the KBHSB solutions and the chosen value of the disk momentum $\ell_{0}$, we found and classified four different possible disk types. Type 1 solutions are characterized by an inner cusp and an outer cusp, between which the disk center lies. Type 2 solutions only exist for prograde disks for the analyzed values of $a$ and are characterized by the absence of an inner cusp, they only possess an outer cusp. The Type 1 solutions can be further subclassified regarding their effective potential values at the cusps. For Type 1a solutions, the effective potential has a higher value at the inner cusp than at the outer cusp. This causes the equi-potential surfaces of the inner cusp to be open towards the center and towards the outer disk regions. The equi-potential surface of the outer cusp fully encloses the disk center region and is closed towards the KBHSB center but open towards the outer region. In the case of Type 1b solutions, the effective potential of the outer cusp has a higher value than the inner cusp. This causes the equi-potential surface of the outer cusp to be open towards the center and towards the outer region. The inner cusp potential surface now fully encloses the disk center, it is closed towards the outer region and open towards the KBHSB center. In the special case of Type 1c, the effective potential has the same value at the inner and outer cusp, the potential surface is fully closed between the inner and outer cusp, and it encloses the disk center but is open towards the KBHSB center and the outer disk region. All mentioned solutions have a high complexity of disk dynamics due to the different conceivable accretion mechanisms regarding the inner as well as the outer cusp. However, direct accretion onto the KBHSB, with the absence of inner-physical disk effects, is only possible for Type 1 solutions, as Type 2 do not possess an inner cusp and are closed towards the KBHSB center.
@@ Acknowledgements.
We would like to thank Betti Hartmann and Jutta Kunz for many of the discussions that led to this paper. R.C. is grateful to CAPES for financial support under Grant No: 88887.371717/2019-00, and would like to thank the University of Oldenburg for hospitality.